TODD Candle Co. produces scented candles. Their most recent income statement is as follows:
All Amounts in Thousands of US $

Revenue $2,400
Variable cost @
$3 per candle 900
Fixed costs 600
Depreciation 400
EBIT 500
Income taxes @ 40% 200
Net Income $300
=====
What is the contribution margin per unit?
What is Accounting Breakeven expressed in number of units sold?
What is the degree of operating leverage?
Answer #1

Number of units sold = 900/3 = 300

Contribution margin per unit = (Sales revenue - Variable costs) / sales

= (2400-900)/300

= $5.

Break even units = Fixed cost /Contribution margin per unit

= 600 / 5

= 120 units

Degree of operating leverage = Contribution margin/EBIT

= (2400-900)/500

= 3

COMMENT IF YOU FACE ANY ISSUES